M&M Milk Chocolate candies come in 7 different colors.  Mars, the parent company of M&Ms used the...

  1. M&M Milk Chocolate candies come in 7 different colors.  Mars, the parent company of M&Ms used the internet to solicit global opinions for a seventh color.  There were 3 colors and this is how Japan voted: 38% chose pink, 36% chose teal and 26% chose purple.  If we pick 3 respondents at random, what is the probability at least one chose pink?
Answer #1

Here' the answer to the question. Please let me know in case you've doubts.

p(pink) = .38

p(not pink) = 1-.38 = .62

n = 3

The above can be treated as params of binomial distribution.

Hence, if we pick 3 respondents at random the possibility at least one chose pink is:

P(X>=1) = P(X=1,2,3)

= P(X=1) + P(X=2) + P(X=3)

= 3C1 * .38^1 *.62^2 + 3C2* .38^2 *.62^1 + 3C3* .38^3 *.62^0

= 0.7617

So, probability at least one choose pink is 0.7617